# Information Density Bounds for Privacy

### v1

@@ Appendix E Proof of Proposition 3
If $P_{Y\mid X}$ satisfies $\varepsilon$-LDP, then it holds that
$\frac{P_{Y\mid X=x}(y)}{P_{Y\mid X=x^{\prime}}(y)}\leq e^{\varepsilon},\quad x,x^{\prime}\in\mathcal{X},\;y\in\mathcal{Y}.$
Fix $x\in\mathcal{X}$ and $y\in\mathcal{Y}$. We can write
$\frac{P_{Y}(y)}{P_{Y\mid X=x}(y)}$ $=\frac{\sum_{x^{\prime}\in\mathcal{X}}P_{Y\mid X=x^{\prime}}(y)P_{X}(x^{\prime})}{P_{Y\mid X=x}(y)}$
$\leq\frac{P_{Y\mid X=x}(y)P_{X}(x)+e^{\varepsilon}P_{Y\mid X=x}(y)(1-P_{X}(x))}{P_{Y\mid X=x}(y)}$
@@
$\Lambda(X\to y)=\log\;\max_{x\in\mathcal{X}}\frac{P_{Y}(y)}{P_{Y\mid X=x}(y)}\leq\log\;\Big{(}e^{\varepsilon}-p_{\mathrm{min}}(e^{\varepsilon}-1)\Big{)}.$
@@ Appendix F Proof of Theorem 4
  1. 1.
@@
    $1=\sum_{x\in\mathcal{X}}P_{X\mid Y=y}(x)$ $=P_{X\mid Y=y}(x^{*})+\sum_{x\in\mathcal{X}\setminus\{x^{*}\}}P_{X\mid Y=y}(x)$
@@
    $\Lambda(X\to y)=\log\;\max_{x\in\mathcal{X}}\frac{P_{X}(x)}{P_{X\mid Y=y}(x)}\leq\log\;\max_{x\in\mathcal{X}}\frac{P_{X}(x)}{1-e^{\varepsilon_{u}}(1-P_{X}(x))}=\log\frac{p_{\mathrm{min}}}{1-e^{\varepsilon_{u}}(1-p_{\mathrm{min}})},$ (71)
@@
    $1=\sum_{x\in\mathcal{X}}P_{X\mid Y=y}(x)$ $=P_{X\mid Y=y}(x^{*})+\sum_{x\in\mathcal{X}\setminus\{x^{*}\}}P_{X\mid Y=y}(x)$
@@
    $\ell(X\to y)=\log\;\max_{x\in\mathcal{X}}\frac{P_{X\mid Y=y}(x)}{P_{X}(x)}\leq\log\;\max_{x\in\mathcal{X}}\frac{1-e^{-\varepsilon_{l}}(1-P_{X}(x))}{P_{X}(x)}=\log\frac{1-e^{-\varepsilon_{l}}(1-p_{\mathrm{min}})}{p_{\mathrm{min}}},$ (73)
@@
Appendix G Proof of Theorem 5
To prove the theorem, we show that if $\operatorname*{ess\,sup}_{P_{Y}}\ell(X\to Y)=\infty$, then $\operatorname*{ess\,sup}_{P_{Y}}\Lambda(X\to Y)=\infty$ as well. To do this, first, we describe the implications of having an unbounded PML. Note that we can write
$\operatorname*{ess\,sup}_{P_{Y}}\ell(X\to Y)$ $=D_{\infty}(P_{XY}\|P_{X}\times P_{Y})=\log\sup_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\;\frac{\operatorname*{ess\,sup}\limits_{P_{X}}P_{Y\mid X}(\mathcal{E}\mid X)}{P_{Y}(\mathcal{E})}.$
Hence, assuming that $P_{XY}\ll P_{X}\times P_{Y}$, $\ell(X\to Y)$ is unbounded if and only if one of the following two conditions hold: either
  1. 1.
    there exists a decreasing sequence of events $\mathcal{A}_{1},\mathcal{A}_{2},\ldots\in\mathcal{S}_{\mathcal{Y}}$ and a set $\mathcal{B}\in\mathcal{S}_{\mathcal{X}}$ with $P_{X}(\mathcal{B})>0$ such that $P_{Y}(\mathcal{A}_{i})\to 0$ but $P_{Y\mid X=x}(\mathcal{A}_{i})\not\to 0$ for all $x\in\mathcal{B}$, or
  2. 2.
    there exists a decreasing sequence of events $\mathcal{A}_{1},\mathcal{A}_{2},\ldots\in\mathcal{S}_{\mathcal{Y}}$ and a set $\mathcal{B}\in\mathcal{S}_{\mathcal{X}}$ with $P_{X}(\mathcal{B})>0$ such that $P_{Y}(\mathcal{A}_{i})\to 0$, and $P_{Y\mid X=x}(\mathcal{A}_{i})\to 0$, but $P_{Y\mid X=x}(\mathcal{A}_{i})$ converges at a strictly slower rate compared to $P_{Y}(\mathcal{A}_{i})$ for all $x\in\mathcal{B}$.
Next, we argue that the first case can never happen when $P_{XY}\ll P_{X}\times P_{Y}$. Recall that $P_{XY}\ll P_{X}\times P_{Y}$ implies that $P_{Y\mid X=x}\ll P_{Y}$ almost surely. Fix $x\in\mathcal{X}$ and let $f=\frac{dP_{Y\mid X=x}}{dP_{Y}}$. Let $(\mathcal{A}_{i})_{i\in\mathbb{N}}$ be a sequence of events with $\lim_{i\to\infty}P_{Y}(\mathcal{A}_{i})=0$. Observe that
$\lim_{i\to\infty}P_{Y\mid X=x}(\mathcal{A}_{i})=\lim_{i\to\infty}\int_{\mathcal{Y}}\mathbf{1}_{\mathcal{A}_{i}}(y)\;P_{Y\mid X=x}(dy)=\lim_{i\to\infty}\int_{\mathcal{Y}}\mathbf{1}_{\mathcal{A}_{i}}(y)f(y)\;P_{Y}(dy).$ (74)
When $f$ is bounded $\lim_{i\to\infty}P_{Y}(\mathcal{A}_{i})=0$ clearly implies that $\lim_{i\to\infty}P_{Y\mid X=x}(\mathcal{A}_{i})=0$. Furthermore, the argument carries over to arbitrary $P_{Y}$-integrable $f$ by first writing $f$ as the limit of a sequence of bounded functions, and then using the dominated convergence theorem.
We conclude that $\ell(X\to Y)$ is unbounded if and only if the second condition above holds. Observe that if there exists a decreasing sequence of events $(\mathcal{A}_{i})_{i\in\mathbb{N}}$ and a non-negligible set $\mathcal{B}\in\mathcal{S}_{\mathcal{X}}$ such that $P_{Y\mid X=x}(\mathcal{A}_{i})$ converges to zero more slowly than $P_{Y}(\mathcal{A}_{i})$ for all $x\in\mathcal{B}$, then there exists a non-negligible set $\mathcal{C}\in\mathcal{S}_{\mathcal{X}}$ such that $P_{Y\mid X=x}(\mathcal{A}_{i})$ converges to zero faster than $P_{Y}({\mathcal{A}_{i}})$ for all $x\in\mathcal{C}$. Thus, we can write
$\operatorname*{ess\,sup}_{P_{Y}}\Lambda(X\to Y)$ $=D_{\infty}(P_{X}\times P_{Y}\|P_{XY})$ (75a)
$=\log\sup_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\;\frac{P_{Y}(\mathcal{E})}{\operatorname*{ess\,inf}\limits_{P_{X}}\;P_{Y\mid X}\mathcal{E}\mid X)}$ (75b)
$\geq\log\lim_{i\to\infty}\;\frac{P_{Y}(\mathcal{A}_{i})}{\operatorname*{ess\,inf}\limits_{P_{X}}\;P_{Y\mid X}(\mathcal{A}_{i}\mid X)}$ (75c)
$\geq\log\lim_{i\to\infty}\;\frac{P_{Y}(\mathcal{A}_{i})}{\operatorname*{ess\,sup}\limits_{P_{X}}\;\Big{(}\mathbf{1}_{\mathcal{C}}(X)\,P_{Y\mid X}(\mathcal{A}_{i}\mid X)\Big{)}}$ (75d)
==$\infty$,= ∞ , (75e)
where (75d) is because the essential infimum of a function is at least as small as its essential supremum over a non-negligible set, and (75e) follows from the definition of set $\mathcal{C}$.
@@ Appendix H Proof of Theorem 6
We show that
$\operatorname*{ess\,sup}_{P_{Y}}\;\Lambda(X\to Y)=\log\sup_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\;\frac{P_{Y}(\mathcal{E})}{\operatorname*{ess\,inf}\limits_{P_{X}}P_{Y\mid X}(\mathcal{E}\mid X)}$ (76)
is finite if and only if
$\log\;\operatorname*{ess\,sup}_{P_{X}\times P_{X}}\;\sup_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\;\frac{P_{Y\mid X}(\mathcal{E}\mid X)}{P_{Y\mid X}(\mathcal{E}\mid X^{\prime})}=\log\sup_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\;\frac{\operatorname*{ess\,sup}\limits_{P_{X}}\;P_{Y\mid X}(\mathcal{E}\mid X)}{\operatorname*{ess\,inf}\limits_{P_{X}}\;P_{Y\mid X}(\mathcal{E}\mid X^{\prime})},$ (77)
is finite. Since $P_{Y}(\mathcal{E})=\mathbb{E}_{X\sim P_{X}}[P_{Y\mid X}(\mathcal{E}\mid X)]\leq\operatorname*{ess\,sup}_{P_{X}}P_{Y\mid X}(\mathcal{E}\mid X)$ for all $\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}$, then the boundedness of (77) implies the boundedness of (76). To prove the other direction, note that if
$\operatorname*{ess\,sup}_{P_{X}\times P_{X}}\;\sup_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\;\frac{P_{Y\mid X}(\mathcal{E}\mid X)}{P_{Y\mid X}(\mathcal{E}\mid X^{\prime})}=\infty,$
then
$(P_{X}\times P_{X})\left(\left\{\sup_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\;\frac{P_{Y\mid X}(\mathcal{E}\mid X)}{P_{Y\mid X}(\mathcal{E}\mid X^{\prime})}>M\right\}\right)>0,$ (78)
for arbitrarily large $M\geq 0$. Put differently, it holds that [^7] If $(P_{X}\times P_{X})\left(\left\{\sup\limits_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\;\frac{P_{Y\mid X}(\mathcal{E}\mid X)}{P_{Y\mid X}(\mathcal{E}\mid X^{\prime})}=\infty\right\}\right)=0$, then there exists $A\geq 0$ such that $\frac{P_{Y\mid X}(\mathcal{E}\mid X)}{P_{Y\mid X}(\mathcal{E}\mid X^{\prime})}\leq A$ almost surely for all $\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}$. This contradicts (78).
$(P_{X}\times P_{X})\left(\left\{\sup_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\;\frac{P_{Y\mid X}(\mathcal{E}\mid X)}{P_{Y\mid X}(\mathcal{E}\mid X^{\prime})}=\infty\right\}\right)>0.$
This, in turn, implies that there exist sets $\mathcal{B},\mathcal{C}\in\mathcal{S}_{\mathcal{X}}$ with $P_{X}(\mathcal{B})P_{X}(\mathcal{C})>0$ such that
$\sup_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\frac{P_{Y\mid X=x}(\mathcal{E})}{P_{Y\mid X=x^{\prime}}(\mathcal{E})}=\infty,$
for all $x\in\mathcal{B}$ and $x^{\prime}\in\mathcal{C}$. Hence, (77) is infinite if and only if one of the following two conditions hold: either
  1. 1.
    there exists an event $\mathcal{A}\in\mathcal{S}_{\mathcal{Y}}$ such that $P_{Y\mid X=x}(\mathcal{A})>0$ for all $x\in\mathcal{B}$ and $P_{Y\mid X=x^{\prime}}(\mathcal{A})=0$ for all $x^{\prime}\in\mathcal{C}$, or
  2. 2.
    there exists a decreasing sequence of events $\mathcal{A}_{1},\mathcal{A}_{2},\ldots\in\mathcal{S}_{\mathcal{Y}}$ such that $P_{Y\mid X=x}(\mathcal{A}_{i})\to 0$ for all $x\in\mathcal{B}$ and $P_{Y\mid X=x^{\prime}}(\mathcal{A}_{i})\to 0$ for all $x^{\prime}\in\mathcal{C}$ as $i\to\infty$, but the convergence on $\mathcal{B}$ is at a strictly slower rate compared to the convergence on $\mathcal{C}$.
Next, we argue that PMC is unbounded under both conditions. Under the first condition, we have
$P_{Y}(\mathcal{A})=\mathbb{E}\Big{[}P_{Y\mid X}(\mathcal{A}\mid X)\Big{]}\geq\mathbb{E}\Big{[}\mathbf{1}_{B}(X)P_{Y\mid X}(\mathcal{A}\mid X)\Big{]}>0,$
and
$\operatorname*{ess\,inf}_{P_{X}}\;P_{Y\mid X}(\mathcal{A}\mid X^{\prime})=0.$
Thus, we get
$\operatorname*{ess\,sup}_{P_{Y}}\Lambda(X\to Y)\geq\log\;\frac{P_{Y}(\mathcal{A})}{\operatorname*{ess\,inf}\limits_{P_{X}}\;P_{Y\mid X}(\mathcal{A}\mid X^{\prime})}=\infty.$
Next, let $\mathcal{S}_{\mathcal{B}}\coloneqq\{\mathcal{F}\cap\mathcal{B}:\mathcal{F}\in\mathcal{S}_{\mathcal{X}}\}$ denote the trace of $\mathcal{S}_{\mathcal{X}}$ on $\mathcal{B}$. Let the measure $Q(\mathcal{F})\coloneqq P_{X}(\mathcal{F})$ for $\mathcal{F}\in\mathcal{S}_{\mathcal{B}}$ denote the restriction of $P_{X}$ to $\mathcal{B}$. If the second condition is satisfied, then we can write
$\operatorname*{ess\,sup}_{P_{Y}}\Lambda(X\to Y)$ $=\log\sup_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\;\frac{P_{Y}(\mathcal{E})}{\operatorname*{ess\,inf}\limits_{P_{X}}\;P_{Y\mid X}(\mathcal{E}\mid X^{\prime})}$ (79a)
$\geq\log\lim_{i\to\infty}\;\frac{P_{Y}(\mathcal{A}_{i})}{\operatorname*{ess\,inf}\limits_{P_{X}}\;P_{Y\mid X}(\mathcal{A}_{i}\mid X^{\prime})}$ (79b)
$=\log\lim_{i\to\infty}\;\frac{\int_{\mathcal{X}}P_{Y\mid X=x}(\mathcal{A}_{i})\,P_{X}(dx)}{\operatorname*{ess\,inf}\limits_{P_{X}}\;P_{Y\mid X}(\mathcal{A}_{i}\mid X^{\prime})}$ (79c)
$\geq\log\lim_{i\to\infty}\;\frac{\int_{\mathcal{B}}P_{Y\mid X=x}(\mathcal{A}_{i})\,P_{X}(dx)}{\operatorname*{ess\,inf}\limits_{P_{X}}\;P_{Y\mid X}(\mathcal{A}_{i}\mid X^{\prime})}$ (79d)
$\geq\log\lim_{i\to\infty}\;\frac{P_{X}(\mathcal{B})\left(\operatorname*{ess\,inf}\limits_{Q}P_{Y\mid X}(\mathcal{A}_{i}\mid X)\right)}{\operatorname*{ess\,inf}\limits_{P_{X}}\;P_{Y\mid X}(\mathcal{A}_{i}\mid X^{\prime})}$ (79e)
$\geq\log\lim_{i\to\infty}\;\frac{P_{X}(\mathcal{B})\left(\operatorname*{ess\,inf}\limits_{Q}P_{Y\mid X}(\mathcal{A}_{i}\mid X)\right)}{\operatorname*{ess\,sup}\limits_{P_{X}}\;\Big{(}\mathbf{1}_{\mathcal{C}}(X^{\prime})\,P_{Y\mid X}(\mathcal{A}_{i}\mid X^{\prime})\Big{)}}$ (79f)
==$\infty$,= ∞ , (79g)
where (79f) is because the essential infimum of a function is at least as small as its essential supremum over a non-negligible set, and (79g) follows by assumption. We conclude that PMC is essentially bounded if and only if it satisfies $\varepsilon$-LDP for a finite value of $\varepsilon$.

### v2

@@ Appendix E Proof of Proposition 3
Let $f_{Y\mid X}=\frac{dP_{Y\mid X}}{d\nu}$ and $f_{Y}=\frac{dP_{Y}}{d\nu}$, where $\nu$ is a dominating $\sigma$-finite measure on $(\mathcal{Y},\mathcal{S}_{\mathcal{Y}})$. If $P_{Y\mid X}$ satisfies $\varepsilon$-LDP, then for all pairs $x,x^{\prime}\in\mathcal{X}$ there exists a measurable set $\mathcal{A}_{x,x^{\prime}}\in\mathcal{Y}$ with probability $P_{Y}(\mathcal{A}_{x,x^{\prime}})=1$ such that
$\frac{f_{Y\mid X=x}(y)}{f_{Y\mid X=x^{\prime}}(y)}\leq e^{\varepsilon},\quad y\in\mathcal{A}_{x,x^{\prime}}.$
Since $\mathcal{X}$ is finite, we can take $\mathcal{A}=\bigcap_{x,x^{\prime}\in\mathcal{X}}\mathcal{A}_{x,x^{\prime}}$ which satisfies $P_{Y}(\mathcal{A})=1$. On this set, we have
$\frac{f_{Y\mid X=x}(y)}{f_{Y\mid X=x^{\prime}}(y)}\leq e^{\varepsilon},\quad x,x^{\prime}\in\mathcal{X},\;y\in\mathcal{A}.$
Fix $x\in\mathcal{X}$ and $y\in\mathcal{A}$. We can write
$\frac{f_{Y}(y)}{f_{Y\mid X=x}(y)}=\frac{\sum_{x^{\prime}\in\mathcal{X}}f_{Y\mid X=x^{\prime}}(y)P_{X}(x^{\prime})}{f_{Y\mid X=x}(y)}$
$\leq\frac{f_{Y\mid X=x}(y)P_{X}(x)+e^{\varepsilon}f_{Y\mid X=x}(y)(1-P_{X}(x))}{f_{Y\mid X=x}(y)}$
@@
$\Lambda(X\to y)$ $=\log\;\max_{x\in\mathcal{X}}\frac{f_{Y}(y)}{f_{Y\mid X=x}(y)}$
$\leq\log\;\Big(e^{\varepsilon}-p_{\mathrm{min}}(e^{\varepsilon}-1)\Big),\quad y\in\mathcal{A}.$
Therefore, the mechanism satisfies $\log\big(e^{\varepsilon}-p_{\mathrm{min}}(e^{\varepsilon}-1)\big)$-PMC.
@@ Appendix F Proof of Theorem 5
Let $f_{Y\mid X}=\frac{dP_{Y\mid X}}{d\nu}$ and $f_{Y}=\frac{dP_{Y}}{d\nu}$. Fix $y\in\mathcal{Y}$. Since $\mathcal{X}$ is finite, we have
$\sup_{P_{X}\in\mathcal{P}_{\mathcal{X}}}\Lambda(X\to y)=\sup_{P_{X}\in\mathcal{P}_{\mathcal{X}}}\log\,\max_{x\in\mathcal{X}}\frac{f_{Y}(y)}{f_{Y\mid X=x}(y)}$
$=\log\,\max_{x\in\mathcal{X}}\frac{\sup\limits_{P_{X}\in\mathcal{P}_{\mathcal{X}}}\sum_{x^{\prime}\in\mathcal{X}}f_{Y\mid X=x^{\prime}}(y)P_{X}(x^{\prime})}{f_{Y\mid X=x}(y)}$
$=\log\,\max_{x\in\mathcal{X}}\frac{\max_{x^{\prime}\in\mathcal{X}}f_{Y\mid X=x^{\prime}}(y)}{f_{Y\mid X=x}(y)}$
$=\log\,\max_{x,x^{\prime}\in\mathcal{X}}\frac{f_{Y\mid X=x^{\prime}}(y)}{f_{Y\mid X=x}(y)}.$
Therefore, for all $\varepsilon\geq 0$ and all $y\in\mathcal{Y}$, we have
$\sup_{P_{X}\in\mathcal{P}_{\mathcal{X}}}\Lambda(X\to y)\leq\varepsilon\iff\log\,\max_{x,x^{\prime}\in\mathcal{X}}\frac{f_{Y\mid X=x^{\prime}}(y)}{f_{Y\mid X=x}(y)}\leq\varepsilon.$
Appendix G Proof of Theorem 6
  1. (i)
@@ Appendix G Proof of Theorem 6
    $1$ $=\sum_{x\in\mathcal{X}}P_{X\mid Y=y}(x)$
    $=P_{X\mid Y=y}(x^{*})+\sum_{x\in\mathcal{X}\setminus\{x^{*}\}}P_{X\mid Y=y}(x)$
@@
    $\begin{split}\Lambda(X\to y)&=\log\;\max_{x\in\mathcal{X}}\frac{P_{X}(x)}{P_{X\mid Y=y}(x)}\\ &\leq\log\;\max_{x\in\mathcal{X}}\frac{P_{X}(x)}{1-e^{\varepsilon_{u}}(1-P_{X}(x))}\\ &=\log\frac{p_{\mathrm{min}}}{1-e^{\varepsilon_{u}}(1-p_{\mathrm{min}})},\end{split}$ (84)
@@
    $1$ $=\sum_{x\in\mathcal{X}}P_{X\mid Y=y}(x)$
    $=P_{X\mid Y=y}(x^{*})+\sum_{x\in\mathcal{X}\setminus\{x^{*}\}}P_{X\mid Y=y}(x)$
@@
    $\ell(X\to y)$ $=\log\;\max_{x\in\mathcal{X}}\frac{P_{X\mid Y=y}(x)}{P_{X}(x)}$
    $\leq\log\;\max_{x\in\mathcal{X}}\frac{1-e^{-\varepsilon_{l}}(1-P_{X}(x))}{P_{X}(x)}$
    $=\log\frac{1-e^{-\varepsilon_{l}}(1-p_{\mathrm{min}})}{p_{\mathrm{min}}},$
@@ Appendix H Proof of Theorem 7
Let $f_{Y\mid X}=\frac{dP_{Y\mid X}}{d\nu}$ and $f_{Y}=\frac{dP_{Y}}{d\nu}$, where $\nu$ is a dominating $\sigma$-finite measure on $(\mathcal{Y},\mathcal{S}_{\mathcal{Y}})$. Recall that by Definition 6, a mechanism satisfies $\varepsilon$-LDP if
$\operatorname*{ess\,sup}_{P_{X}\times P_{X}}\;D_{\infty}\Big(P_{Y\mid X}(\cdot\mid X)\big\|P_{Y\mid X}(\cdot\mid X^{\prime})\Big)$
$=\operatorname*{ess\,sup}_{P_{X}\times P_{X}\times P_{Y}}\,\log\,\frac{f_{Y\mid X}(Y\mid X)}{f_{Y\mid X}(Y\mid X^{\prime})}\leq\varepsilon.$
Note that in the expression for the Rényi divergence, we have taken the essential supremum with respect to $P_{Y}$ since our assumptions in Section II-A ensure that $P_{Y\mid X=x}\ll P_{Y}$ almost surely. Equivalently, there exists a set $\mathcal{E}\in\mathcal{S}_{\mathcal{X}}\otimes\mathcal{S}_{\mathcal{X}}\otimes\mathcal{S}_{\mathcal{Y}}$ with probability one such that
$\frac{f_{Y\mid X}(y\mid x)}{f_{Y\mid X}(y\mid x^{\prime})}\leq e^{\varepsilon},\quad\text{ for all }(x,x^{\prime},y)\in\mathcal{E}.$
For each $(x,y)$ define the section $\mathcal{E}_{x,y}=\{x^{\prime}\in\mathcal{X}:(x,x^{\prime},y)\in\mathcal{E}\}$ and note that $P_{X}(\mathcal{E}_{x,y})=1$ for $P_{X}\times P_{Y}$-a.s. $(x,y)$. Fix such a pair $(x,y)$. Then, we have
$f_{Y}(y)$ $=\int_{\mathcal{X}}f_{Y\mid X}(y\mid x^{\prime})\,P_{X}(dx^{\prime})$
$\leq\int_{\mathcal{E}_{x,y}}f_{Y\mid X}(y\mid x^{\prime})\,P_{X}(dx^{\prime})$
$\leq e^{\varepsilon}f_{Y\mid X}(y\mid x)\int_{\mathcal{E}_{x,y}}P_{X}(dx^{\prime})$
$=e^{\varepsilon}f_{Y\mid X}(y\mid x)P_{X}(\mathcal{E}_{x,y})=e^{\varepsilon}f_{Y\mid X}(y\mid x).$
Hence,
$\frac{f_{Y}(y)}{f_{Y\mid X}(y\mid x)}\leq e^{\varepsilon},\quad P_{X}\times P_{Y}\text{-a.s.},$
i.e., the mechanism satisfies $\varepsilon$-PMC. Applying the same argument, we get $f_{Y}(y)\geq e^{-\varepsilon}f_{Y\mid X}(y\mid x)$ almost surely, so the mechanism satisfies $\varepsilon$-PML.
Conversely, suppose the mechanism satisfies $\varepsilon_{1}$-PMC and $\varepsilon_{2}$-PML. Let
$\mathcal{A}_{1}=\Big\{(x,y):\frac{f_{Y}(y)}{f_{Y\mid X}(y\mid x)}\leq e^{\varepsilon_{1}}\Big\},$
$\mathcal{A}_{2}=\Big\{(x^{\prime},y):\frac{f_{Y\mid X}(y\mid x^{\prime})}{f_{Y}(y)}\leq e^{\varepsilon_{2}}\Big\},$
and note that $(P_{X}\times P_{Y})(\mathcal{A}_{1})=(P_{X}\times P_{Y})(\mathcal{A}_{2})=1$. Define
$\mathcal{A}=\Big\{(x,x^{\prime},y):(x,y)\in\mathcal{A}_{1},\text{ and }(x^{\prime},y)\in\mathcal{A}_{2}\Big\},$
so $\mathcal{A}$ has probability one under the probability measure $P_{X}\times P_{X}\times P_{Y}$ (since the complements of both $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ have probability zero). Moreover, for all $(x,x^{\prime},y)\in\mathcal{A}$, we have
$\frac{f_{Y}(y)}{f_{Y\mid X}(y\mid x)}\leq e^{\varepsilon_{1}},\quad\text{and}\quad\frac{f_{Y\mid X}(y\mid x^{\prime})}{f_{Y}(y)}\leq e^{\varepsilon_{2}},$
and multiplying these inequalities yields
$\frac{f_{Y\mid X}(y\mid x^{\prime})}{f_{Y\mid X}(y\mid x)}\leq e^{\varepsilon_{1}+\varepsilon_{2}},\quad\text{ for all }(x,x^{\prime},y)\in\mathcal{A}.$
Therefore, the mechanism satisfies $\varepsilon_{1}+\varepsilon_{2}$-LDP.